How do I figure out the answer for 40 3/4 - 15 7/8
Answer by richwmiller(17219)

convert 3/4 to eighths
3/4=6/8
+40 6/8
-15 7/8
doesn't work
borrow 8/8 from 40 and add to 6/8
+39 14/8
-15 7/8=
+24 7/8